OpenStudy (anonymous):

If (3,1) is a point on the graph of y=f(x), which of the following must be on the graph of y=1/2f(x)? a. (3,1/2) b. (3/2,1) c. (3/2,1/2) d.(3,1) Please help???

OpenStudy (anonymous):

The answer is A. Because when you have y=cf(x) you multiply the c with x so it would look like this: y=f(c*x). In this case your c=1/2 and your x=1 so, 1*.5= .5 or 1/2

OpenStudy (anonymous):

That makes sense, thank you!